# Stable rank and intrinsic dimensionI.C.F. Ipsen and A.K. Saibaba

### v1

@@ 3.2.2 Adding a matrix of rank 1
The second inequality in (4) means that adding a Hermitian positive semi-definite matrix of rank 1 can increase the intrinsic dimension by at most one.
@@
$\operatorname{\mathrm{intdim}}(\boldsymbol{A}+\boldsymbol{B})-\operatorname{\mathrm{intdim}}(\boldsymbol{A})\leq 1.$
@@
Applying Weyl’s monotonicity theorem (1), $\|\boldsymbol{A}+\boldsymbol{B}\|_{2}\geq\|\boldsymbol{A}\|_{2}$ and $\|\boldsymbol{A}+\boldsymbol{B}\|_{2}\geq\|\boldsymbol{B}\|_{2}$ gives
@@
The following example illustrates that adding a Hermitian positive semi-definite matrix of rank 1 can lower the intrinsic dimension by more than one.
@@
$\operatorname{\mathrm{intdim}}(\boldsymbol{A}+\boldsymbol{B})-\operatorname{\mathrm{intdim}}(\boldsymbol{A})<-1.$
@@ 3.3 Rank of a product
If $\boldsymbol{A}\in{\mathbb{C}}^{m\times n}$ and $\boldsymbol{B}\in\mathbb{C}^{n\times k}$, then [5, Section 0.4.5(c)]
@@
$\operatorname{\mathrm{sr}}(\boldsymbol{A}\boldsymbol{B})\not\leq\operatorname{\mathrm{sr}}(\boldsymbol{A}),\qquad\operatorname{\mathrm{sr}}(\boldsymbol{A}\boldsymbol{B})\not\leq\operatorname{\mathrm{sr}}(\boldsymbol{B}).$
@@
$\operatorname{\mathrm{sr}}(\boldsymbol{A}\boldsymbol{B})$ $\not\leq\min\{\operatorname{\mathrm{sr}}(\boldsymbol{A}),\operatorname{\mathrm{sr}}(\boldsymbol{B})\}$
$\operatorname{\mathrm{sr}}(\boldsymbol{A}\boldsymbol{B})$ $\not\leq\max\{\operatorname{\mathrm{sr}}(\boldsymbol{A}),\operatorname{\mathrm{sr}}(\boldsymbol{B})\}.$
@@
$\operatorname{\mathrm{intdim}}(\boldsymbol{A}\boldsymbol{B})\not\leq\operatorname{\mathrm{intdim}}(\boldsymbol{A}),\qquad\operatorname{\mathrm{intdim}}(\boldsymbol{A}\boldsymbol{B})\not\leq\operatorname{\mathrm{intdim}}(\boldsymbol{B}).$
@@
$\operatorname{\mathrm{intdim}}(\boldsymbol{A}\boldsymbol{B})$ $\not\leq\min\{\operatorname{\mathrm{intdim}}(\boldsymbol{A}),\operatorname{\mathrm{intdim}}(\boldsymbol{B})\}$
$\operatorname{\mathrm{intdim}}(\boldsymbol{A}\boldsymbol{B})$ $\not\leq\max\{\operatorname{\mathrm{intdim}}(\boldsymbol{A}),\operatorname{\mathrm{intdim}}(\boldsymbol{B})\}.$
The stable rank satisfies (5) in the special case where $\boldsymbol{A}$ or $\boldsymbol{B}$ is unitary, or real orthogonal. The intrinsic dimension satisfies (5) in the special case where $\boldsymbol{A}$ or $\boldsymbol{B}$ equals $\alpha\boldsymbol{I}_{n}$ with $\alpha>0$.
@@ 4.1 Multiplication by nonsingular matrices
Multiplication by nonsingular matrices does not change the rank. That is, let $\boldsymbol{A}\in{\mathbb{C}}^{m\times n}$. If $\boldsymbol{B}\in{\mathbb{C}}^{m\times m}$ and $\boldsymbol{C}\in{\mathbb{C}}^{n\times n}$ are nonsingular, then [5, Section 0.4.6(b)]
@@
4.2 Rank of a cross product
The rank of the cross product is equal to the rank of the original matrix. That is, if $\boldsymbol{A}\in{\mathbb{C}}^{m\times n}$ then [5, Section 0.4.6.(d)]
@@ 4.2 Rank of a cross product
Neither the stable rank nor the intrinsic dimension satisfy (7). They are bounded by but not equal to the stable rank or intrinsic dimension of the cross product.
@@
The stable rank and intrinsic dimension do not satisfy (7), that is, the stable rank and intrinsic dimension of a cross product can be strictly smaller than those of the matrix.
@@ 5.1 Review of Schatten p-norms
The $p$-stable rank is defined in terms of Schatten $p$-norms which, in turn, are defined on the singular values of real and complex matrices, and thus special cases of symmetric gauge functions [1, Chapter IV], [7, Section 1], [4, Sections 3.4-3.5], [10].
@@
Let $\boldsymbol{A}\in{\mathbb{C}}^{m\times n}$ have singular values $\sigma_{1}(\boldsymbol{A})\geq\cdots\geq\sigma_{r}(\boldsymbol{A})$, where $r\equiv\min\{m,n\}$. For $1\leq p\leq\infty$, the Schatten $p$-norms are
@@
The Schatten $p$-norms satisfy various properties, including monotonicity, submultiplicativity, strong submultiplicativity, and unitary invariance (symmetric norm).
@@
$\interleave\boldsymbol{A}\boldsymbol{B}\interleave_{p}$ $\leq\interleave\boldsymbol{A}\interleave_{p}\interleave\boldsymbol{B}\interleave_{p}$
@@ 5.2 The p-stable rank
The following generalizes the stable rank relation from Section 2.3 between the cross product and the matrix.
@@
Although the functions in Definition 5.1 are only quasi-norms for $10\leq p<$ 1 rather than norms, because they do not satisfy the triangle inequality [1, Problem IV.5.1], we can extend Definition 5.3 to include this range. Indeed, for $p=0$, we have $\operatorname{\mathrm{sr}}_{0}(\boldsymbol{A})=r$.
@@ 5.3 Sum inequalities for the pth root of the p-stable rank
We show that the $p$th root of the $p$-stable rank satisfies the rank-sum inequality (3), provided both matrices are Hermitian positive semi-definite.
@@
In particular, the square root of the 2-stable rank satisfies the rank-sum inequality (3), provided both matrices are Hermitian positive semi-definite.
@@ 5.3.1 Adding a matrix of rank 1
$p]{\operatorname{\mathrm{sr}}_{p}(\boldsymbol{A}+\boldsymbol{B})}-\sqrt[p]{\operatorname{\mathrm{sr}}_{p}(\boldsymbol{A})}\leq 1.$
@@
Applying Weyl’s monotonicity theorem (1), $\|\boldsymbol{A}+\boldsymbol{B}\|_{2}\geq\|\boldsymbol{A}\|_{2}$ and $\|\boldsymbol{A}+\boldsymbol{B}\|_{2}\geq\|\boldsymbol{B}\|_{2}$ gives
@@
In particular, the square root of the 2-stable rank satisfies the second inequality in (4) for Hermitian positive semi-definite matrices.
@@
$\sqrt{\operatorname{\mathrm{sr}}(\boldsymbol{A}+\boldsymbol{B})}-\sqrt{\operatorname{\mathrm{sr}}(\boldsymbol{A})}\leq 1.$
Theorem 3.5 showed that the 1-stable rank satisfies the second inequality in (4) for Hermitian positive semi-definite matrices.
@@ 5.4 Product inequalities for the p-stable rank
Unfortunately, we were not able to exploit the many existing product inequalities for Schatten $p$-norms [1, Section IV.2], [4, Section 3.5] because they tend to hold only for the Hermitian positive semi-definite polar factor of the product, rather than the product itself.
@@
to the $p$-stable rank of a product gives the upper bound
@@
Below is an extension of Theorem 4.7 that shows that the rank of the cross product is bounded by the rank of the matrix.
@@ 5.5 Conditioning of the pth root of the p-stable rank
Unlike the rank of a matrix, which is ill-posed, the $p$-stable ranks are well-posed, due to their continuous dependence on the matrix elements. In particular, the square root of the stable rank and the intrinsic dimension are well-posed.
@@
Proof 5.22.
If $\operatorname{\mathrm{sr}}_{p}(\boldsymbol{A})=0$ or $\operatorname{\mathrm{sr}}_{p}(\boldsymbol{E})=0$, then the bounds clearly hold. Thus, we assume $\operatorname{\mathrm{sr}}_{p}(\boldsymbol{A})\neq 0$ and $\operatorname{\mathrm{sr}}_{p}(\boldsymbol{E})\neq 0$.
@@
Below is the special case of 2-stable rank.
@@
Below is the special case of 1-stable rank for Hermitian positive semi-definite matrices
@@ Acknowledgement
We thank Srinivas Eswar for helpful discussions.

### v2

@@ 3.2.2 Adding a matrix of rank 1
The second inequality in (8) implies that adding a Hermitian positive semi-definite matrix of rank 1 can increase the intrinsic dimension by at most one.
@@
$1\displaystyle 1\leq\operatorname{\mathrm{intdim}}(\boldsymbol{A}+\boldsymbol{B})\leq\operatorname{\mathrm{intdim}}(\boldsymbol{A})+$ 1 .
@@
Since $\boldsymbol{A}$ and $\boldsymbol{B}$ are Hermitian positive semi-definite, so is $\boldsymbol{A}+\boldsymbol{B}$. In addition, $\operatorname{\mathrm{rank}}(\boldsymbol{B})=1$ implies $\operatorname{\mathrm{rank}}(\boldsymbol{A}+\boldsymbol{B})\geq\operatorname{\mathrm{rank}}(\boldsymbol{B})=1$. Thus $\operatorname{\mathrm{intdim}}(\boldsymbol{A}+\boldsymbol{B})\geq 1$. This establishes the lower bound.
If $\boldsymbol{A}=\boldsymbol{0}$ then $\operatorname{\mathrm{intdim}}(\boldsymbol{A})=0$ and $\operatorname{\mathrm{intdim}}(\boldsymbol{A}+\boldsymbol{B})=\operatorname{\mathrm{intdim}}(\boldsymbol{B})=1$, so that the upper bound holds with equality.
Now assume that $\boldsymbol{A}\neq\boldsymbol{0}$. Applying Weyl’s monotonicity theorem (3), $\|\boldsymbol{A}+\boldsymbol{B}\|_{2}\geq\|\boldsymbol{A}\|_{2}$ and $\|\boldsymbol{A}+\boldsymbol{B}\|_{2}\geq\|\boldsymbol{B}\|_{2}$ gives
@@
If $\boldsymbol{A}=\boldsymbol{0}$, then the lower and upper bounds in Theorem 3.5 hold with equality. The following example illustrates that adding a Hermitian positive semi-definite matrix of rank 1 can lower the intrinsic dimension by more than one, and that for $\boldsymbol{A}\neq\boldsymbol{0}$ the intrinsic dimension of the sum in Theorem 3.5 can be arbitrarily close to one.
@@
$\operatorname{\mathrm{intdim}}(\boldsymbol{A}+\boldsymbol{B})<\operatorname{\mathrm{intdim}}(\boldsymbol{A})-1.$
Furthermore, $\operatorname{\mathrm{intdim}}(\boldsymbol{A}+\boldsymbol{B})\rightarrow 1$ as $\beta\rightarrow\infty$.
@@ 3.3 Rank of a product
If $\boldsymbol{A}\in{\mathbb{C}}^{m\times n}$ and $\boldsymbol{B}\in\mathbb{C}^{n\times k}$, then [11, Section 0.4.5(c)]
@@
$\operatorname{\mathrm{sr}}(\boldsymbol{A}\boldsymbol{B})>\operatorname{\mathrm{sr}}(\boldsymbol{A}),\qquad\operatorname{\mathrm{sr}}(\boldsymbol{A}\boldsymbol{B})>\operatorname{\mathrm{sr}}(\boldsymbol{B}).$
@@
$\operatorname{\mathrm{sr}}(\boldsymbol{A}\boldsymbol{B})$ $>\max\{\operatorname{\mathrm{sr}}(\boldsymbol{A}),\operatorname{\mathrm{sr}}(\boldsymbol{B})\}.$
@@
$\operatorname{\mathrm{intdim}}(\boldsymbol{A}\boldsymbol{B})>\operatorname{\mathrm{intdim}}(\boldsymbol{A}),\qquad\operatorname{\mathrm{intdim}}(\boldsymbol{A}\boldsymbol{B})>\operatorname{\mathrm{intdim}}(\boldsymbol{B}).$
@@
$\operatorname{\mathrm{intdim}}(\boldsymbol{A}\boldsymbol{B})$ $>\max\{\operatorname{\mathrm{intdim}}(\boldsymbol{A}),\operatorname{\mathrm{intdim}}(\boldsymbol{B})\}.$
The stable rank satisfies (9) in the special case where $\boldsymbol{A}$ or $\boldsymbol{B}$ is unitary, or real orthogonal. The intrinsic dimension satisfies (9) in the special case where $\boldsymbol{A}$ or $\boldsymbol{B}$ equals $\alpha\boldsymbol{I}_{n}$ with $\alpha>0$. In general, for $\operatorname{\mathrm{intdim}}(\boldsymbol{A}\boldsymbol{B})$ to be defined, the product $\boldsymbol{A}\boldsymbol{B}$ must also be Hermitian positive definite, which is the case when $\boldsymbol{A}$ and $\boldsymbol{B}$ are simultaneously diagonalizable.
@@ 4.1 Multiplication by nonsingular matrices
Multiplication by nonsingular matrices does not change the rank. That is, let $\boldsymbol{A}\in{\mathbb{C}}^{m\times n}$. If $\boldsymbol{B}\in{\mathbb{C}}^{m\times m}$ and $\boldsymbol{C}\in{\mathbb{C}}^{n\times n}$ are nonsingular, then [11, Section 0.4.6(b)]
@@
4.2 Rank of a cross product matrix
The rank of a cross product matrix is equal to the rank of the original matrix. That is, if $\boldsymbol{A}\in{\mathbb{C}}^{m\times n}$ then [11, Section 0.4.6.(d)]
@@ 4.2 Rank of a cross product matrix
Neither the stable rank nor the intrinsic dimension satisfy (11). They are bounded by but not equal to the stable rank or intrinsic dimension of the cross product matrix.
@@
The stable rank and intrinsic dimension do not satisfy (11), that is, the stable rank and intrinsic dimension of a cross product matrix can be strictly smaller than those of the matrix.
@@ 5.1 Review of Schatten p-norms
The $p$-stable rank is defined in terms of Schatten $p$-norms which, in turn, are defined on the singular values of real and complex matrices, and thus special cases of symmetric gauge functions [2, Chapter IV], [13, Section 1], [10, Sections 3.4-3.5], [20].
@@
Let $\boldsymbol{A}\in{\mathbb{C}}^{m\times n}$ have singular values $\sigma_{1}(\boldsymbol{A})\geq\cdots\geq\sigma_{r}(\boldsymbol{A})$, where $r\equiv\min\{m,n\}$. For $1\leq p\leq\infty$, the Schatten $p$-norms are defined as
@@
The Schatten $p$-norms satisfy various properties, including monotonicity, submultiplicativity, strong submultiplicativity, and unitary invariance (symmetric norm), as shown below.
@@
(14) $\interleave\boldsymbol{A}\boldsymbol{B}\interleave_{p}$ $\leq\interleave\boldsymbol{A}\interleave_{p}\interleave\boldsymbol{B}\interleave_{p}$
@@ 5.2 The p-stable rank
In analogy to (4), the relation (12) between Schatten $p$-norms and two-norm can be improved to
(17) $\interleave\boldsymbol{A}\interleave_{p}$ $p]{\operatorname{\mathrm{sr}}_{p}(\boldsymbol{A})}\>\interleave\boldsymbol{A}\interleave_{\infty}=\sqrt[p]{\operatorname{\mathrm{sr}}_{p}(\boldsymbol{A})}\>\|\boldsymbol{A}\|_{2}.$
Extending the submultiplicative property (5) to general Schatten $p$-norms gives
(18) $\interleave\boldsymbol{A}\boldsymbol{B}\interleave_{p}$ $p]{\min\{\operatorname{\mathrm{sr}}_{p}(\boldsymbol{A}),\operatorname{\mathrm{sr}}_{p}(\boldsymbol{B}),\operatorname{\mathrm{sr}}_{p}(\boldsymbol{A}\boldsymbol{B})\}}\>\interleave\boldsymbol{A}\interleave_{\infty}\interleave\boldsymbol{B}\interleave_{\infty}$
(19) =$p]{\min\{\operatorname{\mathrm{sr}}_{p}(\boldsymbol{A}),\operatorname{\mathrm{sr}}_{p}(\boldsymbol{B}),\operatorname{\mathrm{sr}}_{p}(\boldsymbol{A}\boldsymbol{B})\}}\>\|\boldsymbol{A}\|_{2}\|\boldsymbol{B}\|_{2}.$
Note the difference to (14), which bounds a Schatten $p$-norm in terms of the very same Schatten $p$-norm, rather than the two-norm.
@@
The following generalizes the stable rank relation from Section 2.3 between a matrix and its cross product matrix.
@@
Although the functions in Definition 5.1 are only quasi-norms for $10\leq p<$ 1 rather than norms, because they do not satisfy the triangle inequality [2, Problem IV.5.1], we can extend Definition 5.3 to include this range. Indeed, for $p=0$, we have $\operatorname{\mathrm{sr}}_{0}(\boldsymbol{A})=r$.
@@ 5.3 Sum inequalities for the pth root of the p-stable rank
We show that the $p$th root of the $p$-stable rank satisfies the rank-sum inequality (7), provided both matrices are non-zero Hermitian positive semi-definite.
@@
In particular, the square root of the 2-stable rank satisfies the rank-sum inequality (7), provided both matrices are non-zero Hermitian positive semi-definite.
@@ 5.3.1 Adding a matrix of rank 1
[…]
@@
The proof is similar to that of Theorem 3.5.
Since $\boldsymbol{A}$ and $\boldsymbol{B}$ are Hermitian positive semidefinite, so is $\boldsymbol{A}+\boldsymbol{B}$. In addition, $\operatorname{\mathrm{rank}}(\boldsymbol{B})=1$ implies $\operatorname{\mathrm{rank}}(\boldsymbol{A}+\boldsymbol{B})\geq\operatorname{\mathrm{rank}}(\boldsymbol{B})=1$. Thus $\operatorname{\mathrm{sr}}_{p}(\boldsymbol{A}+\boldsymbol{B})\geq 1$. This establishes the lower bound.
If $\boldsymbol{A}=\boldsymbol{0}$ then $\operatorname{\mathrm{sr}}_{p}(\boldsymbol{A})=0$ and $\operatorname{\mathrm{sr}}_{p}(\boldsymbol{A}+\boldsymbol{B})=\operatorname{\mathrm{sr}}_{p}(\boldsymbol{B})=1$, so that the upper bound holds with equality.
Now assume that $\boldsymbol{A}\neq\boldsymbol{0}$. Applying Weyl’s monotonicity theorem (3), $\|\boldsymbol{A}+\boldsymbol{B}\|_{2}\geq\|\boldsymbol{A}\|_{2}$ and $\|\boldsymbol{A}+\boldsymbol{B}\|_{2}\geq\|\boldsymbol{B}\|_{2}$ gives
@@
If $\boldsymbol{A}=\boldsymbol{0}$, then the lower and upper bounds in Theorem 5.13 hold with equality. The following example illustrates that for $\boldsymbol{A}\neq\boldsymbol{0}$ the $p$-stable rank of the sum in Theorem 5.13 can be arbitrarily close to one.
Example 5.15.
Let
$\boldsymbol{A}=\begin{bmatrix}0&0\\ 0&1\end{bmatrix},\qquad\boldsymbol{B}=\begin{bmatrix}\beta&0\\ 0&0\end{bmatrix},\qquad\boldsymbol{A}+\boldsymbol{B}=\begin{bmatrix}1&0\\ 0&\beta\end{bmatrix}$
where $\beta>1$ where $\operatorname{\mathrm{sr}}_{p}(\boldsymbol{B})=1=\operatorname{\mathrm{rank}}(\boldsymbol{B})$. Then
$\operatorname{\mathrm{sr}}_{p}(\boldsymbol{A}+\boldsymbol{B})=\frac{1^{p}+\beta^{p}}{\beta^{p}}=1+\frac{1}{\beta^{p}}\rightarrow 1\quad\mathrm{as}\ \beta\rightarrow\infty.$
As a particular case of Theorem 5.13, the square root of the 2-stable rank satisfies the second inequality in (8) for Hermitian positive semi-definite matrices.
@@
$\sqrt{\operatorname{\mathrm{sr}}(\boldsymbol{A}+\boldsymbol{B})}\leq\sqrt{\operatorname{\mathrm{sr}}(\boldsymbol{A})}+1.$
Theorem 3.5 shows that the 1-stable rank satisfies the second inequality in (8) for Hermitian positive semi-definite matrices.
@@ 5.4 Product inequalities for the p-stable rank
Unfortunately, we were not able to exploit the many existing product inequalities for Schatten $p$-norms [2, Section IV.2], [10, Section 3.5] because they tend to hold only for the Hermitian positive semi-definite polar factor of the product, rather than the product itself.
@@
and the singular value product inequalities [11, 7.3.P16] to the $p$-stable rank of a product gives the upper bound
@@
As in Section 3.3, for $\operatorname{\mathrm{intdim}}(\boldsymbol{A}\boldsymbol{B})$ to be defined, the product $\boldsymbol{A}\boldsymbol{B}$ must also be Hermitian positive definite, which is the case when $\boldsymbol{A}$ and $\boldsymbol{B}$ are simultaneously diagonalizable.
Below is an extension of Theorem 4.7 that shows that the rank of the cross product matrix is bounded by the rank of the matrix.
@@ 5.5 Conditioning of the pth root of the p-stable rank
Unlike the rank, which is ill-posed, the $p$-stable ranks are well-posed, due to their continuous dependence on the matrix elements. In particular, the square root of the stable rank and the intrinsic dimension are well-posed.
@@
Let $\boldsymbol{A},\boldsymbol{E}\in{\mathbb{C}}^{m\times n}$ with $\boldsymbol{A}\neq\boldsymbol{0}$, $r\equiv\operatorname{\mathrm{rank}}(\boldsymbol{E})$, and $\epsilon\equiv\frac{\|\boldsymbol{E}\|_{2}}{\|\boldsymbol{A}\|_{2}}<1$. Then
@@
If, in addition, $\boldsymbol{A},\boldsymbol{E}\in{\mathbb{C}}^{n\times n}$ are Hermitian positive semi-definite, then
@@
Proof 5.23.
If $\operatorname{\mathrm{sr}}_{p}(\boldsymbol{E})=0$, then the bounds clearly hold. Thus, we assume $\operatorname{\mathrm{sr}}_{p}(\boldsymbol{E})\neq 0$.
@@
Below is the special case of the 2-stable rank.
@@
Below is the special case of the 1-stable rank for Hermitian positive semi-definite matrices
@@ Acknowledgement
We thank Srinivas Eswar for informative discussions, and the reviewers for helpful suggestions.